OpenStudy (anonymous):

Determine r(t) explicitly from the conditions given below: r’’(t)=2t^2i+〖te〗^tj-√tk, r’(4)=-i+5j+k, and r(0)4i+4j+k Determine r(t) explicitly from the conditions given below: r’’(t)=2t^2i+〖te〗^tj-√tk, r’(4)=-i+5j+k, and r(0)4i+4j+k @Mathematics

OpenStudy (anonymous):

Heres my thinking: integrate r'' to r' then use r'(4)=-i+5j+k to find the constants in r

OpenStudy (anonymous):

Then integrate r' to r and use r(0)4i+4j+k to find constants in r

OpenStudy (anonymous):

the j component is te^t

OpenStudy (anonymous):

The way I would find the constants are the by setting the integrated function component to the component result in the r''(4) and substituting 4 for t in the integrated function component

OpenStudy (anonymous):

Like so: Integrate r’’(t): ∫2t^2 i+te^t j-√t k = 2/3 t^3+c_1i+e^t(t-1)*+c_2j+2/3 t^(3/2)+c_3k

OpenStudy (anonymous):

-Use r’(4)=-i+5j+k to find constants: i: 2/3 (4)^3+c_1=-1 ->128/3+c_1=-1->c_1=-131/3

OpenStudy (anonymous):

The reason i am questioning my methodology is cause im getting ugly numbers for c

OpenStudy (agreene):

your idea sounds fine.